# From Federated Learning to $\mathbb{X}$-Learning: Breaking the Barriers of Decentrality Through Random Walks

### v1

@@ 3.3 Perception-Aware Random Walk Design
3.4 Random Walks with Flexible Memory
@@
Future Research Directions The exploration of $\mathbb{X}$L with a single walker opens up various future directions, discussed below. Connection to Graph Signal Processing (GSP): We suggest exploring node importance metrics for $\mathbb{X}$L inspired by core signal processing methods, such as GSP [47, 48, 49, 50, 51]. GSP incorporates graph topology into node signal representation through eigen-decomposition of the graph Laplacian. Developing a unified metric for node importance in $\mathbb{X}$L via GSP calls for (i) defining appropriate signals for the nodes based on their data, (ii) studying interpretations of eigen-decomposition for $\mathbb{X}$L, and (iii) leveraging GSP-inspired learning mechanisms (e.g., graph neural networks) for $\mathbb{X}$L. ML Convergence under Time-Varying Transition Matrices: Characterizing the ML convergence under time-varying transition matrices is crucial and will involve obtaining ML convergence bounds for convex and non-convex loss functions, considering the impacts of random walk trajectory on its model convergence. Comparison with Traditional Markov Chain Designs: Analyzing how the dynamic design of transition matrices differs from time-homogeneous Markov chains considering traditional metrics (e.g., covering time and mixing time [52, 53]) is interesting. This involves demonstrating ‘how’ and ‘why’ traditional random walk visiting patterns may lead to inferior performance when compared to $\mathbb{X}$L-specific strategies. Resource-Aware $\mathbb{X}$L: To derive the optimal walker trajectory and training durations at nodes, future work can formulate optimization problems with competing objective functions, comprising the walker’s ML performance, communication/traversal latency, and energy consumption, under realistic transition costs (e.g., energy and delay) between nodes.
@@
4 $\mathbb{X}$-Learning with Multiple Random Walkers
@@
4.1 Collaboration Among Walkers
Having multiple random walkers independently exploring the network can be inefficient, motivating us to study interactions between the walkers. Drawing from physics concepts on interacting particles [54, 55, 56, 57, 58] and epidemic spreading [59, 60, 61, 62, 63, 64], we explore cooperation between random walkers through information transfer. Specifically, when two walkers meet at a node, they “collide” and aggregate their models (e.g., via weighted averaging their models based on the number of seen datapoints). To show this, we simulate $\mathbb{X}$L over a caveman graph with $10$ cliques, each containing $100$ nodes, where each clique is dominated by the presence of a single label of CIFAR-10. All walkers select their next nodes to jump uniformly at random; however, every $10$ jumps, each pair of walkers aggregate by visiting a predetermined node. Fig. 5 shows that collisions and information exchange are beneficial. Further, it shows the diminishing reward of the number of deployed walkers. Investigating the factors influencing this diminishing reward is an exciting research direction.
4.2 Attraction/Repulsion Between Multiple Random Walkers
@@
Future Research Directions The above exploration of $\mathbb{X}$L with multiple random walkers opens up a series of future directions: Incorporating Caching: The need for physical collisions between walkers can be reduced by assuming that walkers leave a trace/footprint of their ML models on nodes. This trace gradually fades over time to signify the staleness of the left ML model. If a new walker arrives at a node where a trace is still fresh/reliable, it can aggregate its ML model with this trace, while accounting for the age of the trace. From Single-Task to Multi-Task Learning: The walkers can potentially train different ML tasks, introducing multi-task learning [65, 66, 67, 68, 69] into $\mathbb{X}$L. This opens up exploration of (i) aggregation between the walkers based on their task similarity, (ii) optimal walker-to-task allocation, and (iii) resource optimization (e.g., number of SGD iterations of walkers). Security of Random Walkers and their Byzantine Resilience: We suggest studying how exploiting the parameters of random walkers during node traversal (e.g., via implementing an inference attack) in different network neighborhoods (e.g., areas with high vs. low data quality) can reveal information about the data of their previously visited nodes. Additionally, exploring the walkers’ robustness against Byzantine attackers (e.g., malicious random walkers that intentionally aggregate with legitimate walkers to degrade their performance) is an interesting direction. Redefining the Notion of Centrality and Perception: Incorporating multiple random walkers necessitates reevaluating node centrality and walkers’ perceptions defined in Sec. 3. These revamps should be done to avoid redundancy in exploration by assigning different centrality values to a node across walkers.
5 Conclusion
@@
Acknowledgements
Not applicable
Funding
S. Hosseinaliour was partially funded by NSF 2421761 while working on this paper.
Competing interests
The authors declare no competing interests.
Additional information
Correspondence and requests for materials should be addressed to S. Hosseinalipour.

### v2

@@ III-C Perception-Aware Random Walk Design
III-D Random Walks with Flexible Memory
@@
Future Research Directions The exploration of $\mathbb{X}$L with a single walker opens up various future directions, discussed below. Connection to Graph Signal Processing (GSP): We suggest exploring node importance metrics for $\mathbb{X}$L inspired by core signal processing methods, such as GSP [13]. GSP incorporates graph topology into node signal representation through eigen-decomposition of the graph Laplacian. Developing a unified metric for node importance in $\mathbb{X}$L via GSP calls for (i) defining appropriate signals for the nodes based on their data, (ii) studying interpretations of eigen-decomposition for $\mathbb{X}$L, and (iii) leveraging GSP-inspired learning mechanisms (e.g., graph neural networks) for $\mathbb{X}$L. ML Convergence under Time-Varying Transition Matrices: Characterizing the ML convergence under time-varying transition matrices is crucial and will involve obtaining ML convergence bounds for convex and non-convex loss functions, considering the impacts of random walk trajectory on its model convergence. Comparison with Traditional Markov Chain Designs: Analyzing how the dynamic design of transition matrices differs from time-homogeneous Markov chains considering traditional metrics (e.g., covering time and mixing time) is interesting. This involves demonstrating ‘how’ and ‘why’ traditional random walk visiting patterns may lead to inferior performance when compared to $\mathbb{X}$L-specific strategies. Resource-Aware $\mathbb{X}$L: To derive the optimal walker trajectory and training durations at nodes, future work can formulate optimization problems with competing objective functions, comprising the walker’s ML performance, communication/traversal latency, and energy consumption, under realistic transition costs (e.g., energy and delay) between nodes.
@@
IV $\mathbb{X}$-Learning with Multiple Random Walkers
@@
IV-A Collaboration Among Walkers
Having multiple random walkers independently exploring the network can be inefficient, motivating us to study interactions between the walkers. Drawing from physics concepts on interacting particles [14] and epidemic spreading [15], we explore cooperation between random walkers through information transfer. Specifically, when two walkers meet at a node, they “collide” and aggregate their models (e.g., via weighted averaging their models based on the number of seen datapoints). To show this, we simulate $\mathbb{X}$L over a caveman graph with $10$ cliques, each containing $100$ nodes, where each clique is dominated by the presence of a single label of CIFAR-10. All walkers select their next nodes to jump uniformly at random; however, every $10$ jumps, each pair of walkers aggregate by visiting a predetermined node. Fig. 5 shows that collisions and information exchange are beneficial. Further, it shows the diminishing reward of the number of deployed walkers. Investigating the factors influencing this diminishing reward is an exciting research direction.
IV-B Attraction/Repulsion Between Multiple Random Walkers
@@
Future Research Directions The above exploration of $\mathbb{X}$L with multiple random walkers opens up a series of future directions: Incorporating Caching: The need for physical collisions between walkers can be reduced by assuming that walkers leave a trace/footprint of their ML models on nodes. This trace gradually fades over time to signify the staleness of the left ML model. If a new walker arrives at a node where a trace is still fresh/reliable, it can aggregate its ML model with this trace, while accounting for the age of the trace. From Single-Task to Multi-Task Learning: The walkers can potentially train different ML tasks, introducing multi-task learning into $\mathbb{X}$L. This opens up exploration of (i) aggregation between the walkers based on their task similarity, (ii) optimal walker-to-task allocation, and (iii) resource optimization (e.g., number of SGD iterations of walkers). Security of Random Walkers and their Byzantine Resilience: We suggest studying how exploiting the parameters of random walkers during node traversal (e.g., via implementing an inference attack) in different network neighborhoods (e.g., areas with high vs. low data quality) can reveal information about the data of their previously visited nodes. Additionally, exploring the walkers’ robustness against Byzantine attackers (e.g., malicious random walkers that intentionally aggregate with legitimate walkers to degrade their performance) is an interesting direction. Redefining the Notion of Centrality and Perception: Incorporating multiple random walkers necessitates reevaluating node centrality and walkers’ perceptions defined in Sec. III. These revamps should be done to avoid redundancy in exploration by assigning different centrality values to a node across walkers.
V Conclusion